OpenStudy (anonymous):

Find the volume of the given solid: Enclosed by the cylinders z=x^2 and y=x^2 and the planes z=0 and y=4. If someone could help me set this up I'd apperciate it!

OpenStudy (anonymous):

is it 0?

OpenStudy (anonymous):

Volumes are not 0?

OpenStudy (anonymous):

z=x^2 and y=x^2 are parabolic cylinders... I draw this cyclinders and planes in3d cartesian coord. system but i cant find a volume... :S

OpenStudy (anonymous):

Yea... I cant figure out how to set it up.. confused with the whole two cylinders thing

OpenStudy (anonymous):